# Surface Plasmon Polaritons: Creation Dynamics and Interference of Slow and Fast Propagating SPPs at a Temporal Boundary

### v1

@@ III-A Dynamic SPP Formation
We now introduce a temporal interface and examine the field response versus time (normalized time defined as $\left(\nicefrac{{\nu_{phSPP2ss}}}{{\rho}}\right)t$, where $\nu_{phSPP2ss}=\nicefrac{{\omega_{0}}}{{\mathrm{Re}\left\{q_{SPP2ss}\left(\omega_{0}\right)\right\}}}$ is the phase velocity for the steady state SPP in temporal region 2), at various heights above the interface, to demonstrate dynamic SPP formation. We use the configuration in Table I, where the temporal region 2 plasma parameters are $\omega_{p12}=1.5\omega_{0}$, $\Gamma_{d12}=0.001\omega_{p12}$. In Fig. 4, before the time-boundary we plot the electric field in temporal region 1, $E_{z21}$, which is simply the usual sinusoidal response to a time-harmonic source. Starting at the time-boundary, when the media configuration changes from a homogeneous dielectric space to a dielectric-plasma interface (Table I) we have $E_{z22}$, forming a transient response. For reference, we also plot the sinusoidal steady-state response for the temporal region 2 media configuration, $E_{z22ss}$, which the transient field must tend to (for $t\gg\nicefrac{{\nu_{phSPP2ss}}}{{\rho}}$) if the response is stable. The fields are plotted for different observation point heights, compared to the wavelength $\lambda_{0}$, specified in the caption of Fig. 4. We obtain the fields, using the total Green’s function (the principal Green’s function obtained from (27) plus the scattered Green’s function obtained from (28)) in (38), where the scattered Green’s function is not approximated as its residue; it is obtained from the formal integration. We can see that at a height greater than a wavelength $\lambda_{0}$ the field in air (i.e., coupling to a continuum, as opposed to the SPP, which is a guided mode) dominates the total field response in temporal region 2, however, for heights less than a wavelength $\lambda_{0}$, the SPP starts to contribute to the total field. We can also see that the fields are continuous (the field in temporal region 1 (green) and the field in temporal region 2 (blue), which includes the transient response) at the temporal boundary $t=0$.
@@
III-A1 Transient Period and Time to Steady State
We now consider the effect of different permittivity configurations (a change to the plasma material) in temporal region 2 on the SPP transient period and time to steady state. In this case, we use a source point height of $z^{\prime}=\nicefrac{{\lambda_{0}}}{{15}}$ and we use the configuration in Table I. In Fig. 5, where we have plots similar to the plots in Figs. 2 (the 3D plots) and 4 (the line plots), we can see that as we increase the plasma frequency the plasma permittivity becomes more negative (more like a metal), and the electrons in the plasma can more quickly respond to the excitation screening the field, resulting in a less confined SPP and longer SPP wavelengths. Furthermore, more reflection back into the air region can occur. Therefore, as the plasma permittivity becomes more negative, the transient response becomes faster, the SPP amplitude decreases (as we go from subfigure (a) to (d) in Fig. 5 we can see that the overall amplitude decreases as a result of more radiation into the air region), and steady state is approached much sooner.
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
Refer to captionRefer to caption
@@ III-A1 Transient Period and Time to Steady State
III-B Interference of Slow and Fast Propagating SPPs
@@
We consider two completely separate permittivity configuration scenarios, with an in-plane source-observation separation distance of $\rho=6\lambda_{0}$, where $\lambda_{0}=\nicefrac{{2\pi}}{{k_{0}}}\approx 20$ $\mu$m is vacuum wavelength. The first scenario is (a): a slow propagating SPP in temporal region 1, and a fast propagating SPP in temporal region 2. We then swap the configurations for temporal region 1 and 2, to see the opposite SPP response, where we then have (b): a fast propagating SPP in temporal region 1, and a slow propagating SPP in temporal region 2. Note that (a) and (b) are completely separate scenarios, where, in general, a faster propagating SPP corresponds to the plasma with a larger plasma frequency.
@@
Refer to captionRefer to caption
Refer to captionRefer to caption
@@
We also see interference occurring before this time, after the fast propagating SPP (from temporal region 2) arrival time and before it. This interference is occurring between the forward and backward waves (from the SPPs and direct radiation), which have shifted frequencies relative to the those waves in temporal region 1, that result due to the momentum conservation at the temporal boundary. Up to this point we haven’t discussed this much, however, we can see it occurring in some of the other previous plots, where it is more pronounced, e.g., in Fig. 4. Based on these results we can see that it is possible to enhance the SPP field intensity by interfering slow and fast propagating SPPs.

### v2

@@ III-A Dynamic SPP Formation
We now introduce a temporal interface and examine the field response versus time (normalized time defined as $\left(\nicefrac{{\nu_{phSPP2ss}}}{{\rho}}\right)t$, where $\nu_{phSPP2ss}=\nicefrac{{\omega_{0}}}{{\mathrm{Re}\left\{q_{SPP2ss}\left(\omega_{0}\right)\right\}}}$ is the phase velocity for the steady state SPP [54] in temporal region 2), at various heights above the interface, to demonstrate dynamic SPP formation. We use the configuration in Table I, where the temporal region 2 plasma parameters are $\omega_{p12}=1.5\omega_{0}$, $\Gamma_{d12}=0.001\omega_{p12}$. In Fig. 4, before the time-boundary we plot the electric field in temporal region 1, $E_{z21}$, which is simply the usual sinusoidal response to a time-harmonic source. Starting at the time-boundary, when the media configuration changes from a homogeneous dielectric space to a dielectric-plasma interface (Table I) we have $E_{z22}$, forming a transient response. For reference, we also plot the sinusoidal steady-state response for the temporal region 2 media configuration, $E_{z22ss}$, which the transient field must tend to (for $t\gg\nicefrac{{\nu_{phSPP2ss}}}{{\rho}}$) if the response is stable. In other words, $E_{z22ss}$ is the value for the field response resulting from the temporal region 2 media configuration (the field response for that media configuration on its own (no time boundary, no temporal region 1; i.e., it’s as if the media was always there)), which the transient field must tend to once the system settles down (this provides a further check on the validity of the model and associated results). The phase velocity referenced here is determined as described previously in this section. The fields are plotted for different observation point heights, compared to the wavelength $\lambda_{0}$, specified in the caption of Fig. 4. We obtain the fields, using the total Green’s function (the principal Green’s function obtained from (27) plus the scattered Green’s function obtained from (28)) in (38), where the scattered Green’s function is not approximated as its residue; it is obtained from the formal integration. We can see that at a height greater than a wavelength $\lambda_{0}$ the field in air (i.e., coupling to a continuum, as opposed to the SPP, which is a guided mode) dominates the total field response in temporal region 2, however, for heights less than a wavelength $\lambda_{0}$, the SPP starts to contribute to the total field. We can also see that the fields are continuous (the field in temporal region 1 (green) and the field in temporal region 2 (blue), which includes the transient response) at the temporal boundary $t=0$.
[…]
l0ptb40pt(a) Refer to caption
[…]
l0ptb45pt(b) Refer to caption
[…]
l-2.5ptb90pt(c)Refer to caption
[…]
l-2.5ptb90pt(d)Refer to caption
@@
III-B Transient Period and Time to Steady State
We now consider the effect of different permittivity configurations (a change to the plasma material) in temporal region 2 on the SPP transient period and time to steady state, where a transient period after the temporal boundary occurs due to the causal response of the system, i.e., it takes time for the SPP to form after the medium is suddenly changed since the material response (electrons) cannot change instantaneously due to their mass. In this case, we use a source point height of $z^{\prime}=\nicefrac{{\lambda_{0}}}{{15}}$ and we use the configuration in Table I. In Fig. 5, where we have plots similar to the plots in Figs. 2 (the 3D plots) and 4 (the line plots), we can see that as we increase the plasma frequency the plasma permittivity becomes more negative (more like a metal), and the electrons in the plasma can more quickly respond to the excitation screening the field, resulting in a less confined SPP and longer SPP wavelengths. Furthermore, more reflection back into the air region can occur. Therefore, as the plasma permittivity becomes more negative, the transient response becomes faster, the SPP amplitude decreases (as we go from subfigure (a) to (d) in Fig. 5 we can see that the overall amplitude decreases as a result of more radiation into the air region), and steady state is approached much sooner. We also note that amplitude modulation can be seen occurring due to the interference between the forward and backward waves (from the SPPs and direct radiation) in temporal region 2, which have shifted frequencies relative to the waves in temporal region 1, that result due to the momentum conservation at the temporal boundary.
[…]
l0ptb200pt(a) Refer to captionRefer to caption
[…]
l0ptb200pt(b) Refer to captionRefer to caption
[…]
l0ptb200pt(c) Refer to captionRefer to caption
[…]
l0ptb200pt(d) Refer to captionRefer to caption
@@ III-B Transient Period and Time to Steady State
III-C Interference of Slow and Fast Propagating SPPs
@@
[…]
@@
[…]
l0ptb35pt(a) Refer to captionRefer to caption
[…]
l0ptb50pt(b) Refer to captionRefer to caption
@@
We also see interference occurring before this time, after the fast propagating SPP (from temporal region 2) arrival time and before it. This interference is occurring between the forward and backward waves (from the SPPs and direct radiation), which have shifted frequencies relative to the waves in temporal region 1, that result due to the momentum conservation at the temporal boundary. Up to this point we haven’t discussed this much, however, we can see it occurring in some of the other previous plots, where it is more pronounced, e.g., in Fig. 4. Based on these results we can see that it is possible to enhance the SPP field intensity by interfering slow and fast propagating SPPs.